Gold, which has a density of 19.32 g/cm3, is the most ductile metal and can be pressed into a thin leaf or drawn out into a long fiber. (a) If a sample of gold with a mass of 4.981 g, is pressed into a leaf of 4.178 μm thickness, what is the area of the leaf? (b) If, instead, the gold is drawn out into a cylindrical fiber of radius 2.400 μm, what is the length of the fiber?
Answer #1

Gold density   = 19.32 g/cc

a) mass of gold m = 4.981 g

Volume V = m/ = 4.981 /19.32 = 0.2578 cc

thickness of leaf = 4.178E-6 m = 4.178E-4 cm

mass per sq.cm = 4.178E-4 *19.32 = 8.0719 E-3 g

Total mass of the sample = 4.981 g

Area of the gold leaf = 4.981/8.0719E-3 = 617.08 sq.cm

b) gold wire radius r = 2.4E-6 m= 2.4E-4 cm

volume of gold per m length = (2.4E-4)2 *100 = 1.81E-5 cc

mass of wire per 1m = 1.81E-5 *19.32 = 3.496E-4 gm

length of gold wire from 4.981 gm of gold = 4.981/3.496E-4

= 1.425E+4 m = 14.25 km